OpenStudy (anonymous):

Help please. Prove that for every n >= 1, the number 11^n+1 + 12^2n-1 is divisible by 133.

OpenStudy (anonymous):

i'd probably say use mathematical induction

OpenStudy (anonymous):

n=1\[11^{n+1} + 12^{2n-1}=11^2+12=133\]its true for n=1

OpenStudy (anonymous):

now suppose the statement holds for n, prove it holds for n+1

OpenStudy (anonymous):

are u familiar with mathematical induction?

OpenStudy (anonymous):

i got it now thanks.

OpenStudy (anonymous):

very well, your welcome :)

OpenStudy (anonymous):